[Brahms-l] Shift report 20050310 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Mar 10 2005 - 08:02:37 EST
Shifter leader was Hiro
Shifters were 


1.  Beam for 4.5 hour
2.  13660: minimum bias run for mult calibration
    13661 13662: (T3-T5 not working correctly. see E-log.)
3.  T3-T5 problem.  (See elog).

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 10 08:03:16 2005

This archive was generated by hypermail 2.1.8 : Thu Mar 10 2005 - 08:03:30 EST